a box drops down from a lorry while moving with a speed of v zero on the road with inclination θ°, where the mass of the box and kinetic friction coefficient and me and u, respectively. for the moment that the box slides up and reaches a possible maximum height(L).

(a) what is work done done by friction?

(b)what is work done by gravity?



1
Expert's answer
2021-03-23T11:10:58-0400

The initial kinetic energy of the box:


"E_K=\frac12 mv^2."

This energy is wasted to move up the incline (acquire potential energy) and to overcome friction. Assume the maximum height is H above the horizontal:


"E_P=mgH,\\
W=\mu mgx=\mu mg\frac{H}{\text{sin}\theta}."

According to conservation of energy:


"E_K=E_P+W,\\\space\\
\frac12 mv^2=mgH\bigg(1+\frac{\mu}{\text{sin}\theta}\bigg),\\\space\\
H=\frac{v^2\text{ sin}\theta}{2g(\text{sin}\theta+\mu)}.\\\space\\
W=\frac{\mu mhv^2}{2(\sin\theta+\mu)}."